ONE SIDE OF A TRIANGLE IS HALF THE LONGEST SIDE. THE THIRD SIDE IS 16 FEET LESS THAN THE LONGEST SIDE. THE PERIMETER IS 69. FIND ALL THREE SIDES.

Answer by Cromlix(4381)

Hi there,
Make longest side = x
One side = x/2
Third side = x - 16
Perimeter = x + x/2 + x - 16
x + x/2 + x - 16 = 69
Collect like terms
x + x + x/2 = 69 + 16
2x + x/2 = 85
(4x + x)/2 = 85
Cross multiply
5x = 85 * 2 (* = times)
5x = 170
x = 34
Longest side = 34 feet
One side = 17 feet
Third side = 18 feet
Hope this helps :-)